The population mean and standard deviation are given below. Find the required probability and determine whether the given sample mean would be considered unusual.

For a sample of n=64​, find the probability of a sample mean being less than 24.5 if mμ=25 and σ=1.31
Answer #1

Given that ,

mean = = 25

standard deviation = = 1.31

n = 64

= 25

=  / n = 1.31/ 64=0.16

P( < 24.5) = P[( - ) / < (24.5-25) / 0.16]

= P(z <-3.13 )

Using z table

probability = 0.0009